# Using machine learning to measure evidence of students’ sensemaking in physics courses

### v2

@@ Acknowledgements.
- Wojnowicz et al. [2024] M. T. Wojnowicz, K. Gili, P. Rath, E. Miller, J. Miller, C. Hancock, M. O’Donovan, S. Elkin-Frankston, T. T. Brunyé, and M. C. Hughes, Discovering group dynamics in coordinated time series via hierarchical recurrent switching-state models (2024).
- Lerner et al. [2006] R. M. Lerner, W. Damon, E. Thelen, and L. B. Smith, Dynamic systems theories, in Handbook of Child Psychology. Volume 1, Theoretical Models of Human Development (John Wiley & Sons, 2006) 6th ed., pp. 258–312.
Appendix A Physics Problems & Annotations
@@
Characteristics
(2) Does the student identify the orientation (e.g. left or right)?
(3) Does the student identify movement 1 (e.g. slow down)?
(4) Does the student identify movement 2 (e.g. zero)?
(5) Does the student identify movement 3 (e.g. speed up)?
Relationships
(6) Does the student identify the descriptive relationship (e.g position changes is a pos/neg velocity)?
(7) Does the student identify the descriptive relationship (e.g velocity changes is a pos/neg acceleration)?
@@ 1. Which ball hits the ground first?
Characteristics
(5) Does the student identify object 1 positioning (e.g. down)?
(6) Does the student identify object 2 positioning (e.g. horizontal up)?
(7) Does the student identify object 1 motion (e.g. speeding up)?
(8) Does the student identify object 2 motion (e.g. speeding up)?
Relationships
@@ 1. What is the speed of the dog relative to the girl?
Characteristics
(4) Does the student identify the positioning of the dog (e.g. left)?
(5) Does the student identify the positioning of the belt (e.g. right)?
@@
(11) Does the student identify the mechanistic relationship between motion of the observer and motion of the observed? (e.g. the speed of the observer relative to the observed object in motion causes the observed motion)
@@
(4) Does the student identify the object positioning before applied force (e.g. upwards)?
(5) Does the student identify the object positioning after applied force (e.g. horizontal)?
@@ Appendix B Annotation Tables
We provide an annotation summary. We showcase the number of student explanations per problem, and the number of those explanations being labeled as positive for each specific criteria.
@@ Appendix C Loss Function Definition
Here, is the total number of problems, is the total number of student explanations per problem in the dataset , is the total number of domains, and is the set of all context criteria (indexed by ) in a particular domain for a specific problem . In the binary cross-entropy expression, is the ground truth prediction for the student explanation associated with problem with respect to domain and specific criteria . is the corresponding predictive probability.
@@ Appendix E AUPRC Scores
We show the AUPRC scores for each independent classifier in Fig. 5 and the average AUPRC scores across models in Table 5.
| AUPRC (95% CI) | Runtime | |
| BoW | 0.831 (0.788, 0.870) | 1 min |
| BERTfrozen | 0.881 (0.846, 0.911) | 7 min |
| BERTfinetuned | 0.896 (0.863, 0.927) | 160 min |
@@ Appendix F Confusion Matrices
We present confusion matrices for all three methods. To produce binary predictions, we threshold the predicted probability at a value of 0.5. We observe similar performance here as in the AUROC scores (which take all possible thresholds into account). BERTfinetune has the fewest classification errors overall. However, we see that BERTfrozen makes slightly fewer false positive mistakes (45 compared to 47). For cases where one type of error has a larger cost than another, careful model and threshold selection are important.

### v3

@@ Acknowledgements.
- Brown et al. [2020] T. B. Brown, B. Mann, N. Ryder, M. Subbiah, J. Kaplan, P. Dhariwal, A. Neelakantan, P. Shyam, G. Sastry, A. Askell, S. Agarwal, A. Herbert-Voss, G. Krueger, T. Henighan, R. Child, A. Ramesh, D. M. Ziegler, J. Wu, C. Winter, C. Hesse, M. Chen, E. Sigler, M. Litwin, S. Gray, B. Chess, J. Clark, C. Berner, S. McCandlish, A. Radford, I. Sutskever, and D. Amodei, Language models are few-shot learners (2020), arXiv:2005.14165 .
Appendix A Physics problems with criteria questions
@@
(2) Does the student identify the object orientation (e.g. left or right)?
(3) Does the student identify movement 1 of the object(e.g. slow down)?
(4) Does the student identify movement 2 of the object (e.g. zero)?
(5) Does the student identify movement 3 of the object (e.g. speed up)?
@@ 1. Which ball hits the ground first?
(5) Does the student identify object 1 spatial organization (e.g. down)?
(6) Does the student identify object 2 spatial organization (e.g. horizontal up)?
(7) Does the student identify object 1 motion post the drop (e.g. speeding up)?
(8) Does the student identify object 2 motion post the drop (e.g. speeding up)?
@@ 1. What is the speed of the dog relative to the girl?
(4) Does the student identify the spatial organization of the dog (e.g. left)?
(5) Does the student identify the spatial organization of the belt (e.g. right)?
@@
(11) Does the student identify the mechanistic relationship between motion of the observer and motion of the observed? (e.g. the observer moving relative to the object impacts the observed motion of the object)
@@
(4) Does the student identify the object spatial organization before applied force (e.g. upwards)?
(5) Does the student identify the object spatial organization after applied force (e.g. horizontal)?
@@ Appendix B Annotation tables
We provide an annotation summary. We showcase the number of student responses per problem, and the number of those explanations being labeled as positive for each specific criteria.
@@ Appendix C Loss function definition
Here, is the total number of problems, is the total number of student responses per problem in the dataset , is the total number of domains, and is the set of all context criteria (indexed by ) in a particular domain for a specific problem . In the binary cross-entropy expression, is the ground truth prediction for the student response associated with problem with respect to domain and specific criteria . is the corresponding predictive probability. Note that by ground truth, we mean the human-annotated labels provided in this work.
@@ Appendix E AUPRC Scores
We show the test set AUPRC scores across domains for each classifier in Fig. 4. The micro-averaged test set AUPRC score for each model is in Table 6.
| AUPRC (95% CI) | Train time | |
| BoW | 0.829 (0.786, 0.868) | 1 min |
| BERTfrozen | 0.876 (0.842, 0.906) | 7 min |
| BERTfinetuned | 0.895 (0.862, 0.926) | 160 min |
@@ Appendix F Confusion matrices
Below, we present for each method a confusion matrix covering all yes/no criteria questions for the 15-student test set. To produce binary predictions, we threshold the predicted probability at a value of 0.5. We observe similar performance here as in the AUROC scores (which take all possible thresholds into account). BERTfinetune has the fewest classification errors overall. However, we see that BERTfrozen makes the same number of false positive mistakes (45). For deployment applications where one type of error has a larger cost than another, careful model and threshold selection are important.
BoW on test set:
BERTfrozen on test set:
BERTfinetuned on test set: